А B 10 kg rigid body at A slides down constant velocity when z= 28° what will be the acceleration of the block when z =53°
Answer #1

Answer pen Efy = may since constant Speed Qy=0 Na mg co(280) = 0 N - (10x9-81) CO3(284) = 0 mg 287 . N = 86.6/57 N 70 Q= simi